Mathematics 243 – Algebraic Structures Definitions Quiz 3 – September 21, 2012

Your Name:

A) (5) What does it mean for f : A → A to be a permutation of A?

B) (5) Is f : Z → Z defined by

f(x) = \{ \begin{matrix}x & if x < 0 \\ x − 1 & if x ≥ 0 \end{matrix}

a permutation of Z? Why or why not?